Find the LCM of (100, 14, 20)

Give the principle please

Answer by askmemath(368)

Write the Prime Number factors of each number
100 = 25x4 = 5^2 x 2^2
14 = 2^1 x 7^1
20 = 5x4 = 5^1 x 2^2
Pick the highest number power value of each prime number from all the numbers
So we get 5^2 x 7^1 x 2^2 = 700
And that's your L.C.M or Least Common Multiple